OpenStudy (anonymous):

y = 4x + 9 y = x solve by substitution. please help!

OpenStudy (anonymous):

since y = x you can write x for y in the first equation x = 4x + 9 x - 4x = 9 -3x = 9 x = -3

OpenStudy (anonymous):

how did you get -4 ?

OpenStudy (anonymous):

and -3x ?

OpenStudy (anonymous):

got it , thank you!!!!!!!!!!! :)

OpenStudy (anonymous):

i need help in one more question, if you have time?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

y = -3x - 5 y = 4x + 9

OpenStudy (anonymous):

be right back, but you can show me how to do it, and ill give u a medal when i come back (10 mins) thank you!

OpenStudy (anonymous):

we can calute be substitution method ....4x+9=-3x-5 4x-3x=-5-9 x=-14 now we get x=-14 put the vaue of x we get value of y\].

OpenStudy (anonymous):

@justinee u undrstand

OpenStudy (anonymous):

yes, but why is it 4x - 3x ?